the market for lemons has 10 potential customers each having an individual curve p=101-10Qi where p is price in dollars per cup and Qi is the number of cups demanded per week by the ith consumer. find the market demand curve using algebra.draw an individual demand curve and the market demand curve.what is the quantity demanded by each consumer and the market as a whole when lemon is priced at p=$/cup?


1
Expert's answer
2021-08-06T15:27:06-0400
"10Q_i=101-p"

"Q=\frac{101-p}{10}\times10=101-p"